# The conformal Laplacian and positive scalar curvature metrics on manifolds with boundary

### v4

@@ 3.1. Results on Ric and $A_{g}$
$+e^{-2\phi}\left(-2(n-2)\nabla_{n}\phi\nabla_{n}\phi+2(n-2)\lvert d\phi\rvert_{g}^{2}g_{nn}-(n-2)(n-1)\lvert d\phi\rvert_{g}^{2}\right).$
@@
(13) $\begin{split}&\qquad\left(-2\text{Ric}_{\tilde{g},\tilde{n}\tilde{n}}+R_{\tilde{g}}\right)|_{\partial M}\\ &=e^{-2\phi}\left(-2\text{Ric}_{g,nn}+2(n-2)\left(\nabla_{n}\nabla_{n}\phi-\nabla_{n}\phi\nabla_{n}\phi\right)+\left(2\Delta_{g}\phi+2(n-2)\lvert d\phi\rvert_{g}^{2}\right)g_{nn}\right)\\ &\qquad+e^{-2\phi}\left(R_{g}-2(n-1)\Delta_{g}\phi-(n-1)(n-2)\lvert d\phi\rvert_{g}^{2}\right)\\ &=e^{-2\phi}\left(-2\text{Ric}_{g,nn}+4u^{-1}\nabla_{n}\nabla_{n}u-3(n-2)\nabla_{n}\phi\nabla_{n}\phi\right)\\ &\qquad+e^{-2\phi}\left(u^{-1}\cdot\frac{4}{n-2}\Delta_{g}u+(n-2)\lvert d\phi\rvert_{g}^{2}\right)\cdot g_{nn}\\ &\qquad\qquad+e^{-2\phi}\left(R_{g}-u^{-1}\frac{4(n-1)}{n-2}\cdot\Delta_{g}u\right)\\ &=e^{-2\phi}u^{-1}\cdot\frac{1}{n-2}\left(-(4n-4-4g_{nn})\Delta_{g}u+(4n-8)\nabla_{n}\nabla_{n}u+(n-2)R_{g}u\right)\\ &\qquad+e^{-2\phi}u^{-1}\cdot\frac{1}{n-2}\left(-2(n-2)u\text{Ric}_{g,nn}+(n-2)^{2}u\lvert d\phi\rvert_{g}^{2}g_{nn}-3(n-2)^{2}u\nabla_{n}\phi\nabla_{n}\phi\right).\end{split}$
Our immediate goal is to prove the existence of a smooth solution of the following PDE, which contains the linear terms in the second to the last line of (13), i.e.,
@@
Under our standing assumption $\eta_{1}>0$, we may assume $R_{g}>0$ by Theorem 3.1 (where $\tilde{g}$ in that result is now relabeled $g$). We may also assume $g_{nn}<1$ on $\partial M$ by appropriate choices of local coordinates. In other words, for coordinates $(x^{1},\ldots,x^{n})$ near $p\in\partial M$ with $\nu_{p}=\partial/\partial x^{n}$, we define new coordinates $(y^{1}=x^{1},\ldots,y^{n-1}=x^{n-1},y^{n}=\alpha^{-1}x^{n})$ for a constant $\alpha<1.$ Since $\nu$ is globally defined, in the $y$ coordinates $g_{nn}=\alpha^{2}$ on all of $\partial M.$ (12) and (13) still hold in the $y$ coordinates.
@@
where we assume $g^{in}=0$ for $i\neq n$, and $\xi^{\perp}:=\xi-\xi_{n}dy^{n}.$ In addition, the zeroth order term $(n-2)R_{g}$ is positive, so the differential operator in (14) is injective. Applying the standard Lax-Milgram argument, we conclude that there exists $u\in H^{1}(M,g)$ solving (14). The smoothness of $u$ follows from standard elliptic regularity arguments. ∎
We investigate the relationship between $R_{\tilde{g}}$ on $\partial M$ and $R_{i^{*}\tilde{g}}$ under the conformal change $\tilde{g}=u^{p-2}g$, where $u$ is obtained by first solving (14) on $C$ for an appropriate $F$, and then extending the solution to a positive function $u$ on $M$ by partition of unity. Of course, $u$ may not solve (14) outside the collar region.
We first choose a smooth $F$ on $\partial M$ such that $F-2(n-2){\rm Ric}_{g,nn}>0\;{\rm on}\;\partial M.$ Fix an integer $q\gg 1$ such that the Sobolev embedding $W^{2,q}(C,g)\hookrightarrow\mathcal{C}^{1,\alpha}(C)$ is continuous. We then extend $F$ smoothly to all of $C$ such that $\lVert F\rVert_{\mathcal{L}^{q}(C,g)}\ll 1$, i.e. the $\mathcal{L}^{q}$-norm is small enough, where this size will be determined later. So far, we have
@@
(21) $-(4n-4-4g_{nn})\Delta_{g}u_{1}$ $+(4n-8)\nabla_{n}\nabla_{n}u_{1}+(n-2)R_{g}u_{1}=F+(n-2)(1+\epsilon)R_{g}D\;{\rm in}\;{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}C},$
@@
Taking the conformal transformation $\tilde{g}=u_{1}^{p-2}g:=e^{\phi_{1}}g,$ it follows from (13) and (21) that
$\qquad\left(-2\text{Ric}_{\tilde{g},\tilde{n}\tilde{n}}+R_{\tilde{g}}\right)|_{\partial M}$
@@
$\qquad+e^{-2\phi_{1}}u_{1}^{-1}\cdot\frac{1}{n-2}\left(-2(n-2)u_{1}\text{Ric}_{g,nn}+(n-2)^{2}u_{1}\lvert d\phi_{1}\rvert_{g}^{2}g_{nn}-3(n-2)^{2}u_{1}\nabla_{n}\phi_{1}\nabla_{n}\phi_{1}\right)|_{\partial M}$
@@
We now extend $u_{1}$ from $C$ to $M$ so that $u_{1}$ is positive on all of $M$. $u_{1}$ satisfies (14) near $\partial M$. From the Gauss-Codazzi equation (7), we conclude the following:
Proposition 3.2.
If $(M^{n},\partial M,g)$ has $\eta_{1}>0$ and $\partial M$ totally geodesic, then there exists a metric $\tilde{g}\in[g]$ such that $R_{\imath^{*}\tilde{g}}>0$ and $\partial M$ remains totally geodesic for $\tilde{g}$.
When $\dim M=3$, $R_{\imath^{*}\tilde{g}}>0$ implies $\chi(\partial M_{i})>0$; when $\dim M\geqslant 4$, $R_{\imath^{*}\tilde{g}}>0$ implies $\zeta_{1}>0$. Using Theorem 2.1, we conclude that
@@
If $(M^{n},\partial M,g)$ has $\eta_{1}>0$ and $\partial M$ totally geodesic, then there exists a metric $\tilde{g}$ conformal to $g$ such that $R_{\tilde{g}}>0$ and $R_{\imath^{*}\tilde{g}}>0$.
This result gives a conformal geometric obstruction on $\partial M$ to the positivity of $\Box_{g}$ on $M$.
Proposition 3.3.
@@
    If $n\geqslant 3$ and $R_{\bar{g}}\leqslant 0$ everywhere on $X$, there exists no compact manifold $(M,g)$ with $\eta_{1}>0$ and totally geodesic boundary, such that $\partial M=X$ and $\imath^{*}g$ is conformal to $\bar{g}$.
@@
    If $n\geqslant 3$ and $\zeta_{1}\leqslant 0$ on $X$, there exists no compact manifold $(M,g)$ with $\eta_{1}>0$ and totally geodesic boundary, such that $\partial M=X$ and $\imath^{*}g$ is conformal to $\bar{g}$.
@@
    If $n=2$ and $\chi(X_{i})\leqslant 0$ for any component $X_{i}$ of $X$, there is no compact $3$-manifold $(M,g)$ with $\eta_{1}>0$ and totally geodesic boundary, such that $\partial M=X$ and $\imath^{*}g$ is conformal to $\bar{g}$; similarly, there is no $3$-manifold $(M,g)$ with $\eta_{1}>0$, such that $\partial M=X$, $i^{*}g$ conformal to $\bar{g}$, and $\bar{g}$ has nonpositive Yamabe invariant on any $X_{i}.$
@@
For (i), assume such an $M$ exists. $R_{\bar{g}}\leq 0$ implies $\zeta_{1,\bar{g}}\leq 0.$ By the hypotheses and Prop. 3.2, there exists $\tilde{g}$ conformal to $g$ with $R_{\imath^{*}\tilde{g}}>0.$ Thus $\zeta_{1,\imath^{*}\tilde{g}}>0.$ However, $\imath^{*}\tilde{g}$ is conformal to $\bar{g}$ by hypothesis, so the sign of $\zeta_{1}$ for these metrics must be the same.
@@
For (iii), if $g$ exists, there exists $\tilde{g}$ conformal to $g$ with $R_{\imath^{*}\tilde{g}}>0$ by Prop. 3.2. This implies $\chi(X_{i})>0$ for all $i$. $R_{\imath^{*}\tilde{g}}>0$ also implies positive Yamabe invariant on each $X_{i}.$
@@
To restate Prop. 3.3(i) geometrically, if $R_{\bar{g}}\leqslant 0$ on $X$, then there exists no compact manifold $(M,g)$ with $R_{g}>0,$ $A_{g}=0$, such that $\partial M=X$ and $\imath^{*}g$ is conformal to $\bar{g}$. This is easy if $g$ is a product metric near $\partial M$: $A_{g}$ automatically vanishes and $R_{i^{*}g}=R_{g}>0$ on $\partial M$, so $\zeta_{1,i^{*}g}\geqslant 0.$ On the other hand, $R_{\bar{g}}\leqslant 0$ implies $\zeta_{1,\bar{g}}\leqslant 0.$ As in the previous proof, this is impossible. The amount of analysis in §3 indicates the large gap between the product and nonproduct cases.
@@
[14, Lemma 2.4] assumes that $\partial M$ is totally geodesic and has positive Yamabe invariant, both conditions on $\partial M.$ By Theorem 3.1, the positivity of the Yamabe invariant follows from the assumption $\eta_{1}>0$, an interior condition.
@@
[28, Thm. 1.1] On $(M^{n},\partial M,g)$, there exists a metric $\tilde{g}$ conformal to $g$ with $R_{\tilde{g}}$ constant on the interior of $M$ and $R_{i^{*}\tilde{g}}$ constant on $\partial M.$
When $\eta_{1}>0$, Theorem 3.1 implies that $R_{\tilde{g}}$ is a positive constant. With the help of Theorem 3.2 and Corollary 3.2, we can determine that $R_{i^{*}\tilde{g}}$ is also a positive constant. This is an improvement of Theorem. 2.1 and Corollary. 3.2.
@@
Let $(M^{n},\partial M,g)$ have totally geodesic boundary. If $\eta_{1}>0$, there exists a metric $\tilde{g}\in[g]$ with $R_{\tilde{g}}$ a positive constant on the interior of $M$ and $R_{i^{*}\tilde{g}}$ a positive constant on $\partial M.$
@@
By Theorem 3.2 and Corollary 3.2, $\zeta_{1}>0$ when $n>3$ and $\chi(\partial M_{i})>0$ when $n=3$. The rest of the proof follows exactly as in Theorem 2.1. ∎
@@
Let $(M^{n},\partial M,g)$ have $\partial M$ totally geodesic and spin. If $\eta_{1}>0$, then $(\partial M,i^{*}g)$ has no nontrivial harmonic spinors.
@@ 4. Positivity of the conformal Laplacian: A quasi-topological obstruction
The index $\operatorname{ind}(\not{\partial}_{\tilde{g}})$ of the Dirac operator with APS boundary conditions vanishes by [16, Thm. 4]: If $h_{\tilde{g}}\geqslant 0$, then the smallest eigenvalue $\lambda_{1}$ of ∂̸$\not{\partial}^{E}∂$ with APS boundary conditions satisfies
@@
where the second line uses [16, (1)]: $X\langle\psi,\phi\rangle_{S}=\langle\nabla^{S}_{X}\psi,\phi\rangle+\langle\psi,\nabla_{X}^{S}\phi\rangle.$ Thus we obtain the twisted version of [16, (1)] from the untwisted version.
@@
from the Lichnerowicz formula ∂̸*∂̸$\not{\partial}^{*}\not{\partial}=(\nabla^{S})^{*}\nabla^{S}+\frac{R}{4}.∂$ As a final example, in our notation [16, (9)] is
@@
This is the twisted version of [16, (9)].
A straightforward but long check yields the twisted versions of (1)-(17), cumulating in the twisted version of [16, Thm. 4]: we still have $\lambda_{1}^{2}\geqslant[(n+1)/4n]\cdot\min_{M}R_{\tilde{g}}.$ Thus the index of ∂̸$\not{\partial}^{E}_{\tilde{g}}∂$ vanishes. ∎
@@
The APS theorem for manifolds with non-product metric near the boundary [11, §3] is
@@
where $\hat{A}(\Omega_{g})$ is the $\hat{A}$-polynomial in the curvature of $g$, and $P(g)$ is the Chern-Simons form related to the $\hat{A}$-polynomial. As a polynomial in the Pontrjagin classes, $\hat{A}(\Omega_{g})$ is a conformal invariant. ${\rm Dim\ ker}(\not{\partial}_{i^{*}g})$ is a conformal invariant, since the Dirac operator is conformally covariant, i.e.. for $\tilde{g}=e^{2\phi}g$, we have ∂̸$\not{\partial}_{\tilde{g}}=e^{-(n+1)\phi/2}\not{\partial}_{g}e^{(n-1)\phi/2}∂$ [15, §4], [17, §1.4]. The eta invariant is also a conformal invariant [20]. Finally, for $g_{1}$ a smoothing on $N=\partial M\times[0,1]$ of the metric $g|_{\partial M}$ at $t=0$ to a product metric near $t=1,$ we have $\int_{\partial M}P(g)=\int_{N}\hat{A}(\Omega_{g_{1}}).$ We can clearly construct $\tilde{g}_{1}$ conformal to $g_{1}$ on $N$ in the obvious notation. As above, this implies that $\int_{\partial M}P(g)$ is a conformal invariant. Thus $\operatorname{ind}(\not{\partial}_{g})=\operatorname{ind}(\not{\partial}_{\tilde{g}})$.
@@
(i) Let $(M^{n},\partial M,g)$ be a spin manifold with boundary with $\eta_{1}>0$, and either totally geodesic boundary or $\zeta_{1}>0$. Assume that $\pi_{1}(\partial M)$ admits a unitary representation $\alpha$ that extends to a unitary representation of $\pi_{1}(M).$ Then $\xi_{\alpha}(\partial M,g)=0.$
(ii) If $(M,g)$ has $R_{g}>0$ and totally geodesic boundary, and we have the representation hypothesis, then $\xi_{\alpha}(\partial M,g)=0.$
@@
As usual, (i) implies (ii). For (i), the left hand side of (24) vanishes for $\tilde{g}$ as in Lem. 4.1, and hence for $g$ by Lemma 4.2. $P(g)=P(\Omega_{g},\omega_{g})$ is a an invariant polynomial in the components of the curvature and connection forms of $g$ [11, §3]. On the right hand side, by Corollary 3.4 there is a different metric $g^{\prime}$ conformal to $g$ with ker(∂̸$\not{\partial}_{i^{*}g^{\prime}})=0( ∂$ on $\partial M$. By the conformal invariance of this kernel, ker(∂̸$\not{\partial}_{i^{*}g})=0( ∂$. Therefore,
@@
(ii) (26) is proved in [2, Thm. 3.9], if $M$ admits a psc metric $g$ which is a product near $\partial M.$ Such a metric has vanishing second fundamental form, so $h_{g}=0$. Thus these metrics have $\eta_{1}>0$, and so form a special case of the Theorem.

### v5

@@ 3.2. Positivity with totally umbilic boundary
$\qquad+e^{-2\phi}\left(-2(n-2)\nabla_{n}\phi\nabla_{n}\phi+2(n-2)\lvert d\phi\rvert_{g}^{2}g_{nn}-(n-2)(n-1)\lvert d\phi\rvert_{g}^{2}\right).$
@@
${\left(-2\text{Ric}_{\tilde{g},\tilde{n}\tilde{n}}+R_{\tilde{g}}\right)|_{\partial M}}$
$=e^{-2\phi}\left(-2\text{Ric}_{g,nn}+2(n-2)\left(\nabla_{n}\nabla_{n}\phi-\nabla_{n}\phi\nabla_{n}\phi\right)+\left(2\Delta_{g}\phi+2(n-2)\lvert d\phi\rvert_{g}^{2}\right)g_{nn}\right)$
$\qquad+e^{-2\phi}\left(R_{g}-2(n-1)\Delta_{g}\phi-(n-1)(n-2)\lvert d\phi\rvert_{g}^{2}\right)$
$=e^{-2\phi}\left(-2\text{Ric}_{g,nn}+4u^{-1}\nabla_{n}\nabla_{n}u-3(n-2)\nabla_{n}\phi\nabla_{n}\phi\right)$
(17) $\qquad+e^{-2\phi}\left(u^{-1}\cdot\frac{4}{n-2}\Delta_{g}u+(n-2)\lvert d\phi\rvert_{g}^{2}\right)\cdot g_{nn}$
$\qquad\qquad+e^{-2\phi}\left(R_{g}-u^{-1}\frac{4(n-1)}{n-2}\cdot\Delta_{g}u\right)$
$=e^{-2\phi}u^{-1}\cdot\frac{1}{n-2}\left(-(4n-4-4g_{nn})\Delta_{g}u+(4n-8)\nabla_{n}\nabla_{n}u+(n-2)R_{g}u\right)$
$\qquad+e^{-2\phi}u^{-1}\cdot\frac{1}{n-2}\left(-2(n-2)u\text{Ric}_{g,nn}+(n-2)^{2}u\lvert d\phi\rvert_{g}^{2}g_{nn}-3(n-2)^{2}u\nabla_{n}\phi\nabla_{n}\phi\right).$
The next step is to prove the existence of a smooth solution of the following PDE, which contains the linear terms in the second to the last line of (3.2), i.e.,
@@
Under our standing assumption $\eta_{1}>0$, we may assume $R_{g}>0$ by Theorem 3.1 (where $\tilde{g}$ in that result is now relabeled $g$). We may also assume $g_{nn}<1$ on $\partial M$ by appropriate choices of local coordinates. In other words, for coordinates $(x^{1},\ldots,x^{n})$ near $p\in\partial M$ with $\nu_{p}=\partial/\partial x^{n}$, define new coordinates $(y^{1}=x^{1},\ldots,y^{n-1}=x^{n-1},y^{n}=\alpha^{-1}x^{n})$ for a constant $\alpha<1.$ Since $\nu$ is globally defined, in the $y$ coordinates $g_{nn}=\alpha^{2}$ on all of $\partial M.$ (16) and (3.2) still hold in the $y$ coordinates.
@@
where $\xi^{\perp}:=\xi-\xi_{n}dy^{n},$ and we use $g^{in}=0$ for $i\neq n$. In addition, the zeroth order term $(n-2)R_{g}$ is positive, so the differential operator in (18) is injective. Applying the standard Lax-Milgram argument, we conclude that there exists $u\in H^{1}(M,g)$ solving (18). The smoothness of $u$ follows from standard elliptic regularity arguments. ∎
We now investigate the relationship between $R_{\tilde{g}}$ on $\partial M$ and $R_{i^{*}\tilde{g}}$ under the conformal change $\tilde{g}=u^{p-2}g$, where $u$ is obtained by first solving (18) on $C$ for an appropriate $F$, and then extending the solution to a positive function $u$ on $M$ by partition of unity. Of course, $u$ may not solve (18) outside the collar region.
Choose $F\in C^{\infty}(\partial M)$ such that $F-2(n-2){\rm Ric}_{g,nn}>0\;{\rm on}\;\partial M.$ Fix an integer $q\gg 1$ such that the Sobolev embedding $W^{2,q}(C,g)\hookrightarrow\mathcal{C}^{1,\alpha}(C)$ is continuous. Extend $F$ smoothly to all of $C$ such that $\lVert F\rVert_{\mathcal{L}^{q}(C,g)}\ll 1$, i.e. the $\mathcal{L}^{q}$-norm is small enough, where this size will be determined later. So far, we have
@@
(25) $-(4n-4-4g_{nn})\Delta_{g}u_{1}$ $+(4n-8)\nabla_{n}\nabla_{n}u_{1}+(n-2)R_{g}u_{1}=F+(n-2)(1+\epsilon)R_{g}D\;{\rm in}\;C,$
@@
Taking the conformal transformation $\tilde{g}=u_{1}^{p-2}g:=e^{\phi_{1}}g,$ it follows from (3.2) and (25) that
${\left(-2\text{Ric}_{\tilde{g},\tilde{n}\tilde{n}}+R_{\tilde{g}}\right)|_{\partial M}}$
@@
$\qquad+e^{-2\phi_{1}}u_{1}^{-1}\cdot\frac{1}{n-2}\left(-2(n-2)u_{1}\text{Ric}_{g,nn}+(n-2)^{2}u_{1}\lvert d\phi_{1}\rvert_{g}^{2}g_{nn}\right)$
$\qquad-e^{-2\phi_{1}}u_{1}^{-1}\cdot\frac{1}{n-2}\left(3(n-2)^{2}u_{1}\nabla_{n}\phi_{1}\nabla_{n}\phi_{1}\right)|_{\partial M}$
@@
We now extend $u_{1}$ from $C$ to $M$ so that $u_{1}$ is positive on all of $M$. $u_{1}$ satisfies (18) near $\partial M$. By (15), we conclude the following:
Theorem 3.1.
If $(M^{n},\partial M,g)$ has $\eta_{1}>0$ and $\partial M$ totally umbilic, then there exists a metric $\tilde{g}$ conformal to $g$ such that $R_{\tilde{g}}>0$ and $R_{\imath^{*}\tilde{g}}>0$.
This result gives a conformal geometric obstruction on $\partial M$ to the positivity of $\Box_{g}$ on $M$, and can be thought of as an obstruction to filling.
@@
    If $n\geqslant 3$ and $R_{\bar{g}}\leqslant 0$ everywhere on $X$, there exists no compact manifold $(M,g)$ with $\eta_{1}>0$ and totally umbilic boundary, such that $\partial M=X$ and $\imath^{*}g$ is conformal to $\bar{g}$.
@@
    If $n\geqslant 3$ and $\zeta_{1}\leqslant 0$ on $X$, there exists no compact manifold $(M,g)$ with $\eta_{1}>0$ and totally umbilic boundary, such that $\partial M=X$ and $\imath^{*}g$ is conformal to $\bar{g}$.
@@
    If $n=2$ and $\chi(X_{i})\leqslant 0$ for any component $X_{i}$ of $X$, there is no compact $3$-manifold $(M,g)$ with $\eta_{1}>0$ and totally umbilic boundary, such that $\partial M=X$ and $\imath^{*}g$ is conformal to $\bar{g}$; similarly, there is no $3$-manifold $(M,g)$ with $\eta_{1}>0$, such that $\partial M=X$, $i^{*}g$ conformal to $\bar{g}$, and $\bar{g}$ has nonpositive Yamabe invariant on any $X_{i}.$
@@
For (i), assume such an $M$ exists. $R_{\bar{g}}\leq 0$ implies $\zeta_{1,\bar{g}}\leq 0.$ By the hypotheses and Thm. 3.1, there exists $\tilde{g}$ conformal to $g$ with $R_{\imath^{*}\tilde{g}}>0.$ Thus $\zeta_{1,\imath^{*}\tilde{g}}>0.$ However, $\imath^{*}\tilde{g}$ is conformal to $\bar{g}$ by hypothesis, so the sign of $\zeta_{1}$ for these metrics must be the same.
@@
For (iii), if $g$ exists, there exists $\tilde{g}$ conformal to $g$ with $R_{\imath^{*}\tilde{g}}>0$ by Thm. 3.1. This implies $\chi(X_{i})>0$ for all $i$. $R_{\imath^{*}\tilde{g}}>0$ also implies positive Yamabe invariant on each $X_{i}.$
@@
To restate Cor. 3.2(i) geometrically, if $R_{\bar{g}}\leqslant 0$ on $X$, then there exists no compact manifold $(M,g)$ with $R_{g}>0$ and $\partial M$ totally umbilic, such that $\partial M=X$ and $\imath^{*}g$ is conformal to $\bar{g}$. This is easy if $g$ is a product metric near $\partial M$: $R_{i^{*}g}=R_{g}>0$ on $\partial M$, so $\zeta_{1,i^{*}g}\geqslant 0.$ On the other hand, $R_{\bar{g}}\leqslant 0$ implies $\zeta_{1,\bar{g}}\leqslant 0.$ As in the previous proof, this is impossible. The amount of analysis in §3 indicates the large gap between the product and nonproduct cases.
@@
An interesting contrast to Cor. 3.2(i) is in [19, Lem. 2.4], where they prove that if $(X,\bar{g})$ has positive Yamabe invariant and nonvanishing conformal invariant $I$, then no $M$ exists with a totally geodesic psc metric conformal to $\bar{g}$ on $X$. We obtain the same conclusion (with totally umbilic boundary) under the hypothesis of nonpositive Yamabe invariant.
@@
[33, Thm. 1.1] On $(M^{n},\partial M,g)$, there exists a metric $\tilde{g}$ conformal to $g$ with $R_{\tilde{g}}$ constant on the interior of $M$ and $R_{i^{*}\tilde{g}}$ constant on $\partial M.$
When $\eta_{1}>0$, Thm. 3.2 implies that $R_{\tilde{g}}$ is a positive constant. With the help of Thm. 3.1, we can determine that $R_{i^{*}\tilde{g}}$ is also a positive constant. This is an improvement of Thm. 2.1.
@@
Let $(M^{n},\partial M,g)$ have totally umbilic boundary. If $\eta_{1}>0$, there exists a metric $\tilde{g}\in[g]$ with $R_{\tilde{g}}$ a positive constant on the interior of $M$ and $R_{i^{*}\tilde{g}}$ a positive constant on $\partial M.$
@@
By Thm. 3.1, $\zeta_{1}>0$ when $n>3$ and $\chi(\partial M_{i})>0$ when $n=3$. The rest of the proof follows exactly as in Thm. 2.1 (with $R_{g_{1}}$ a positive constant). ∎
@@
Let $(M^{n},\partial M,g)$ have $\partial M$ totally umbilic and spin. If $\eta_{1}>0$, then $(\partial M,i^{*}g)$ has no nontrivial harmonic spinors.
@@ 4. Positivity of the conformal Laplacian: A quasi-topological obstruction
The index $\operatorname{ind}(\not{\partial}_{\tilde{g}})$ of the Dirac operator with APS boundary conditions vanishes by [21, Thm. 4]: If $h_{\tilde{g}}\geqslant 0$, then the smallest eigenvalue $\lambda_{1}$ of ∂̸$\not{\partial}^{E}∂$ with APS boundary conditions satisfies
@@
where the second line uses [21, (1)]: $X\langle\psi,\phi\rangle_{S}=\langle\nabla^{S}_{X}\psi,\phi\rangle+\langle\psi,\nabla_{X}^{S}\phi\rangle.$ Thus we obtain the twisted version of [21, (1)] from the untwisted version.
@@
from the Lichnerowicz formula ∂̸∂̸$\not{\partial}^{*}\not{\partial}=(\nabla^{S})^{*}\nabla^{S}+\frac{R}{4}.∂$ As a final example, in our notation [21, (9)] is
@@
This is the twisted version of [21, (9)].
A straightforward but long check yields the twisted versions of (1)-(17), cumulating in the twisted version of [21, Thm. 4]: we still have $\lambda_{1}^{2}\geqslant[(n+1)/4n]\cdot\min_{M}R_{\tilde{g}}.$ Thus the index of ∂̸$\not{\partial}^{E}_{\tilde{g}}∂$ vanishes. ∎
@@
The APS theorem for manifolds with non-product metric near the boundary [15, §3] is
@@
where $\hat{A}(\Omega_{g})$ is the $\hat{A}$-polynomial in the curvature of $g$, and $P(g)$ is the Chern-Simons form related to the $\hat{A}$-polynomial. As a polynomial in the Pontrjagin classes, $\hat{A}(\Omega_{g})$ is a conformal invariant. ${\rm Dim\ ker}(\not{\partial}_{i^{*}g})$ is a conformal invariant, since the Dirac operator is conformally covariant, i.e.. for $\tilde{g}=e^{2\phi}g$, we have ∂̸$\not{\partial}_{\tilde{g}}=e^{-(n+1)\phi/2}\not{\partial}_{g}e^{(n-1)\phi/2}∂$ [20, §4], [22, §1.4]. The eta invariant is also a conformal invariant [25]. Finally, for $g_{1}$ a smoothing on $N=\partial M\times[0,1]$ of the metric $g|_{\partial M}$ at $t=0$ to a product metric near $t=1,$ we have $\int_{\partial M}P(g)=\int_{N}\hat{A}(\Omega_{g_{1}}).$ We can clearly construct $\tilde{g}_{1}$ conformal to $g_{1}$ on $N$ in the obvious notation. As above, this implies that $\int_{\partial M}P(g)$ is a conformal invariant. Thus $\operatorname{ind}(\not{\partial}_{g})=\operatorname{ind}(\not{\partial}_{\tilde{g}})$.
@@
(i) Let $(M^{n},\partial M,g)$ be a spin manifold with boundary with $\eta_{1}>0$, and either totally umbilic boundary or $\zeta_{1}>0$. Assume that $\pi_{1}(\partial M)$ admits a unitary representation $\alpha$ that extends to a unitary representation of $\pi_{1}(M).$ Then $\xi_{\alpha}(\partial M,g)=0.$
(ii) If $(M,g)$ has $R_{g}>0$ and totally umbilic boundary, and we have the representation hypothesis, then $\xi_{\alpha}(\partial M,g)=0.$
@@
As usual, (i) implies (ii). For (i), the left hand side of (28) vanishes for $\tilde{g}$ as in Lem. 4.1, and hence for $g$ by Lemma 4.2. $P(g)=P(\Omega_{g},\omega_{g})$ is a an invariant polynomial in the components of the curvature and connection forms of $g$ [15, §3]. On the right hand side, by Corollary 3.4 there is a different metric $g^{\prime}$ conformal to $g$ with ker(∂̸$\not{\partial}_{i^{*}g^{\prime}})=0( ∂$ on $\partial M$. By the conformal invariance of this kernel, ker(∂̸$\not{\partial}_{i^{*}g})=0( ∂$. Therefore,
@@
(ii) (30) is proved in [4, Thm. 3.9], if $M$ admits a psc metric $g$ which is a product near $\partial M.$ Such a metric has vanishing second fundamental form, so $h_{g}=0$. Thus these metrics have $\eta_{1}>0$, and so form a special case of the Theorem.